OpenStudy (anonymous):

medical x-rays, with a wavelength of about 10^-10m, can penetrate completely through your skin. -ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wave length about 1000 times the wavelength of an x-ray. find the wavelength of ultraviolet rays. (need to show work)

OpenStudy (ash2326):

X rays wavelength= \(10^{-10} m\) UV rays have a wavelength 1000 times that of X-ray so UV rays wavelength=\(1000\times 10^{-10} meters\) Can you simplify this?

OpenStudy (anonymous):

\[1\times10^{-3}\times10^{-10}\]

OpenStudy (anonymous):

\[so 1\times10^{-13^{?}}\] ?

OpenStudy (ash2326):

\[10^3\times 10^{-10}=10^{-7}\]

OpenStudy (anonymous):

ok i get it! thank you

OpenStudy (ash2326):

Cool:)